what is the value of y according to the following code

int x= 10, y

x= 2*3

y=x-x/2


1
Expert's answer
2020-11-04T17:54:32-0500

x = 2*3 = 6;

y = x - x/2 = 6 - 6/2 = 6 - 3 = 3;

Answer: y = 3